### v2

@@ 1 Introduction
We evaluated Marconi on a wide range of workloads, request arrival patterns, cache sizes, and Hybrid model architectures. Overall, we find that Marconi improves cache hit rates by an average of 4.5-34.4$\times$ compared to state-of-the-art caching systems (e.g., vLLM Kwon et al. (2023), SGLang Zheng et al. (2023b)) that are extended to support Hybrid models. The win in token hit rate translates to latency savings of 36.1-71.1% (103.3-617.0ms) for P95 TTFT. Microbenchmarks show that Marconi performs better in scenarios with longer contexts, higher ratios of SSM layers, and larger SSM state dimensions — trends that align with recent model developments.
@@ 2.1 State Space Models and Hybrid Models
During prefill, generating the first token depends on all previous tokens. In Transformers, the self-attention mechanism calculates how each token in the sequence “attends” to every other token. Consequently, the Attention mechanism incurs quadratic computational complexity Dao et al. (2022), which quickly bottlenecks GPU compute as sequence lengths scale Agrawal et al. (2024). Furthermore, the size of the KVs in Attention layers grows linearly with sequence length, resulting in a large inference-time memory requirement Wu et al. (2024); Lee et al. (2024); Gao et al. (2024).
@@ 4.1 Cache Admission
Comparison with SGLang. SGLang Zheng et al. (2023b) is an existing prefix caching system for Transformers that also leverages a radix tree for mapping sequences to tokens and their KVs. Different from SGLang, Marconi uses the past requests in the radix tree to determine which SSM states to cache by performing speculative insertions before prefilling an upcoming sequence. Further, Marconi’s philosophy of judicious state admission fundamentally differs from SGLang, which admits the states of all tokens during admission and is no different from other prefix caching systems like vLLM Kwon et al. (2023). We show how these differences yield superior cache efficiency in §5.
@@ 4.2 Cache Eviction
Managing the balance. Marconi manages the balance between favoring recency and FLOP efficiency by tuning $\alpha$. A higher $\alpha$ emphasizes FLOP efficiency, while setting $\alpha$ to 0 falls back to LRU. Marconi manages the balance by observing the workload and retrospectively setting the best configuration. On startup, Marconi sets $\alpha$ to 0 until the first eviction. Afterward, Marconi takes a snapshot of the radix tree and enters a bootstrap period, continuing to use LRU while bookkeeping token-level information of requests during the bootstrapping phase. The bootstrap period includes 5×5$\times$5 × the number of requests seen before the first eviction, capturing a representative workload sample. Once sufficient requests have been observed, Marconi asynchronously launches a grid search over possible $\alpha$ values by replaying the bootstrap requests. This grid search is parallelized across CPU cores, significantly speeding up the tuning process, typically taking just a few seconds, often shorter than the time required to prefill and decode a single request. After the grid search, Marconi adopts the $\alpha$ value that maximizes the hit rate.
@@ 5.2 End-to-End Results
Figure 9: Distribution of P95 TTFT relative to no prefix caching.
@@ 6 Related Work
(Hybrid) recurrent subquadratic models. There has been a resurgence of recurrent/linear models in the recent years: RWKV Peng et al. (2023), RetNet Sun et al. (2023), GLA Yang et al. (2023), Griffin De et al. (2024), RecurrentGemma Botev et al. (2024), xLSTM Beck et al. (2024), Test-Time Training (TTT) Sun et al. (2024), DeltaNet Yang et al. (2024b), B’MOJO Zancato et al. (2024), etc. Importantly, although these models have different state updating rules Yang et al. (2023; 2024b), they all update their (large) model states recurrently. As such, we only evaluate Mamba/SSMs, a representative architecture. The properties we summarize in §3 also apply to these recurrent layers (with slight variations on FLOP and state size), and Marconi can be extended to support prefix caching for all Hybrid models with recurrent layers.

### v3

@@ 1 Introduction
We evaluated Marconi on a wide range of workloads, request arrival patterns, cache sizes, and Hybrid model architectures. Overall, we find that Marconi improves cache hit rates by an average of 4.5-34.4$\times$ compared to state-of-the-art caching systems (e.g., vLLM Kwon et al. (2023), SGLang Zheng et al. (2023b)) that are extended to support Hybrid models. The win in token hit rate translates to latency savings of 36.1-71.1% (103.3-617.0ms) for P95 TTFT. Microbenchmarks show that Marconi performs better in scenarios with longer contexts, higher ratios of SSM layers, and larger SSM state dimensions — trends that align with recent model developments. Marconi is open sourced at https://github.com/ruipeterpan/marconi.
@@ 2.1 State Space Models and Hybrid Models
During prefill, generating the first token depends on all previous tokens. In Transformers, the self-attention mechanism calculates how each token in the sequence “attends” to every other token. Consequently, the Attention mechanism incurs quadratic computational complexity Dao et al. (2022), which quickly bottlenecks GPU compute as sequence lengths scale Agrawal et al. (2024). Furthermore, the size of the KVs in Attention layers grows linearly with sequence length, resulting in a large inference-time memory requirement Wu et al. (2024); Lee et al. (2024).
@@ 4.1 Cache Admission
Comparison with SGLang. SGLang Zheng et al. (2023b) is an existing prefix caching system for Transformers that also leverages a radix tree for mapping sequences to tokens and their KVs. Different from SGLang, Marconi uses the past requests in the radix tree to determine which SSM states to cache by performing speculative insertions before prefilling an upcoming sequence. Further, Marconi’s philosophy of judicious state admission fundamentally differs from SGLang, which admits the states of all tokens during admission and is no different from other prefix caching systems like vLLM Kwon et al. (2023). We show how these differences yield superior cache efficiency in §5. Our system doesn’t invent new data structures for KV cache management, and instead builds on the rich prior work in this space that also aims to manage prefixes’ model states efficiently. Our main contribution lies in being the first system to redesign caching policies and their interactions with these data structures to practically address the unique properties of emerging Hybrid models.
@@ 4.2 Cache Eviction
Managing the balance. Marconi manages the balance between favoring recency and FLOP efficiency by tuning $\alpha$. A higher $\alpha$ emphasizes FLOP efficiency, while setting $\alpha$ to 0 falls back to LRU. Marconi manages the balance by observing the workload and retrospectively setting the best configuration. On startup, Marconi sets $\alpha$ to 0 until the first eviction. Afterward, Marconi takes a snapshot of the radix tree and enters a bootstrap period, continuing to use LRU while bookkeeping token-level information of requests during the bootstrapping phase. The bootstrap period includes $5\times$ 15 × the number of requests seen before the first eviction, capturing a representative workload sample. Once sufficient requests have been observed, Marconi asynchronously launches a grid search over possible $\alpha$ values by replaying the bootstrap requests. This grid search is parallelized across CPU cores, significantly speeding up the tuning process, typically taking just a few seconds, often shorter than the time required to prefill and decode a single request. After the grid search, Marconi adopts the $\alpha$ value that maximizes the hit rate.
Comparisons with existing size-based eviction algorithms. Cost-aware cache eviction for objects with variable sizes is a well-studied problem (e.g., GDSF Cherkasova (1998)). The KV size in Attention layers scales with sequence length, serving as a proxy for compute savings from cache hits, whereas SSM states are fixed-sized irrespective of sequence length or compute savings. Thus, size fails as a proxy in Hybrid LLM inference, where longer sequences (with greater compute savings) are represented by equally sized SSM states. We note that our system and techniques are complementary to such prior schemes; just as we combine FLOP efficiency with recency, our focus is to augment the estimators for cache entry importance used in prior work with a factor tailored specifically for Hybrid model reuse benefits.
@@ 5.2 End-to-End Results
Figure 9: Distribution of P95 TTFT relative to no prefix caching.
@@ 6 Related Work
(Hybrid) recurrent subquadratic models. There has been a resurgence of recurrent/linear models in the recent years: RWKV Peng et al. (2023), RetNet Sun et al. (2023), GLA Yang et al. (2023), Griffin De et al. (2024), RecurrentGemma Botev et al. (2024), xLSTM Beck et al. (2024), Test-Time Training (TTT) Sun et al. (2024), DeltaNet Yang et al. (2024b; c), B’MOJO Zancato et al. (2024), MambaFormer Park et al. (2024), Titans Behrouz et al. (2024), Lightning Attention Li et al. (2025), Hunyuan-TurboS hun (2025), Nemotron-H nem (2025), etc. Importantly, although these models have different state updating rules Yang et al. (2023; 2024c), they all update their (large) model states recurrently. As such, we only evaluate Mamba/SSMs, a representative architecture. The properties we summarize in §3 also apply to these recurrent layers (with slight variations on FLOP and state size), and Marconi can be extended to support prefix caching for all Hybrid models with recurrent layers.
@@ A.1 FLOP Efficiency Analysis
Appendix B Artifact Appendix
B.1 Abstract
In this artifact, we describe how to reproduce all experiments in this paper. The evaluation utilizes request arrival traces from LMSys, ShareGPT, and SWEBench, all tokenized using the meta-llama/Llama-2-7b-hf tokenizer for consistency. Key experiments involve running vLLM+, eviction policies V1 (SGLang+), V2 (Marconi), and V3 (offline-optimal, static-$\alpha$ oracle policy – not included in the paper) across various dataset/arrival rate/cache size combinations. This artifact supports easy customizations, including implementing additional eviction policies and evaluating new datasets and model configurations. For a detailed version of the artifact appendix, please see https://github.com/ruipeterpan/marconi/blob/main/artifact_evaluation.md.
B.2 Artifact check-list (meta-information)
  • Program: Python
  • Disk space required: $\sim$20 GB
  • Time needed to prepare workflow: $\sim$1 hour
  • Time needed to complete experiments: $\sim$12 hours
  • Publicly available: Yes
  • Code licenses: CC BY-NC
  • Archived DOI: 10.5281/zenodo.14970139
B.3 Description
B.3.1 How delivered
The artifact may be downloaded from zenodo at https://zenodo.org/records/14970139 or cloned from the GitHub repository at https://github.com/ruipeterpan/marconi. The request arrival traces (with token-level information) can be downloaded from Google Drive by following the documentation below.
B.3.2 Hardware dependencies
We have tested Marconi on Cloudlab Duplyakin et al. (2019) nodes with Ubuntu 22.04 and Python 3.11.9. Due to the request arrival traces containing token IDs for each request, they require a total disk size of $\sim$7 GB to house.
B.3.3 Software dependencies
We have prepared a conda environment file that lists all dependencies and their versions.
B.3.4 Datasets
To ensure consistency when evaluating the same trace on different model architectures, we use the same tokenizer to produce the request arrival traces. For convenience, we compress the processed traces ($\sim$700M/6.3G pre/post compression) and host them on Google Drive.
B.4 Installation
The software dependencies can be installed via conda using the environment configuration file provided in the artifact. The traces can be downloaded via this link.
B.5 Experiment workflow
The sweep of all experiments (combinations of different cache sizes and request arrival patterns on different datasets) can be done by running bash run_all_experiments.sh, which invokes policy_exploration.py and does the following for each experiment configuration (dataset/arrival rate/cache size combination):
  • $\bullet$
    Runs vLLM+
  • $\bullet$
    Runs eviction policy V1, which represents SGLang+
  • $\bullet$
    Runs eviction policy V2, which represents Marconi
  • $\bullet$
    Runs eviction policy V3, which represents an offline-optimal, static-$\alpha$ oracle policy (the results weren’t included in the paper). This policy sweeps over possible values of $\alpha$ and selects the one that maximizes the hit rate
Running bash run_all_experiments.sh creates three log files in /logs: lmsys.txt, sharegpt.txt, and swebench.txt. Each file contains the output log of all evaluations on the sweep of configurations for this dataset. Once the log files have been generated, plotting scripts can be run to analyze and plot the results.
B.6 Evaluation and expected result
All plotting scripts are under /plotting and can be run once the sweep of experiments on all configurations has finished:
  • $\bullet$
    Fig. 7: token_hit_rate.py
  • $\bullet$
    Fig. 8: sglang_comparison.py
  • $\bullet$
    Fig. 9: ttft.py
  • $\bullet$
    Fig. 10: fine_grained_analysis.py
  • $\bullet$
    Fig. 11: microbenchmark_contention.py
  • $\bullet$
    Fig. 12a: microbenchmark_layer_composition.py
  • $\bullet$
    Fig. 12b: microbenchmark_dstate.py
  • $\bullet$
    Fig. 13a and 13b: microbenchmark_arrivalrate.py
B.7 Experiment customization
  • $\bullet$
    Additional eviction policies can be easily implemented in radix_cache_hybrid.py by adding a new evict_policy_version
  • $\bullet$
    Additional model configurations can be applied by editing the default configurations (NVIDIA’s Attention-Mamba2 7B Hybrid model) in policy_exploration.py
  • $\bullet$
    Additional datasets can be evaluated by producing compatible traces as the ones generated by /utils/generate_trace.py
B.8 Methodology
Submission, reviewing, and badging methodology: